### v3

# Simulating Non-Markovian Open Quantum Dynamics with Neural Quantum States

Abstract
Reducing computational scaling for simulating non-Markovian dissipative dynamics using artificial neural networks is both a major focus and formidable challenge in open quantum systems. To enable neural quantum states (NQSs), we encode environmental memory in dissipatons (quasiparticles with characteristic lifetimes), yielding the dissipaton-embedded quantum master equation (DQME). The resulting NQS-DQME framework achieves a compact representation of both many-body correlations and non-Markovian memory. Benchmarking against numerically exact hierarchical equations of motion confirms NQS-DQME maintains comparable accuracy while enhancing scalability and interpretability. This methodology opens new paths to explore non-Markovian open quantum dynamics in previously intractable systems.
Introduction. Many-body open quantum systems (OQSs) have attracted wide attention due to their profound applications across diversified fields [Engel2007, doi:10.1126/science.1235820, RevModPhys.94.045008]. Accurately characterizing quantum correlations is crucial for the studies of many-body OQSs. For instance, the Kondo correlation between a magnetic molecule (the primary system) and a substrate (the environment) critically determines the electric conductance across the molecule-substrate interface [wilson1975renormalization, doi:10.1126/science.280.5363.567, doi:10.1126/science.1113449, li2020molecular], exemplifying spatially extended quantum correlations. Similarly, non-Markovian memory effects arising from system-environment coupling [breuer2016colloquium, de2017dynamics, PhysRevLett.133.140403] can lead to complex scaling relations in the dissipative dynamics of OQSs [anders2005real, Din24174120], demonstrating temporally persistent quantum correlations [PhysRevLett.111.086601, doi:10.1126/science.aay6779, doi:10.1126/science.abg8223, Ding25084114].
Numerically exact approaches [Tam18030402, Lam193721, PhysRevLett.82.1801, 10.1063/1.1647528, PhysRevLett.123.050601, PhysRevLett.88.170407, suess2014hierarchy, PhysRevLett.115.266802, PhysRevLett.130.186301, PhysRevB.87.115115, PhysRevLett.88.256403, RevModPhys.92.011001, 10.1063/1.3173823, RevModPhys.93.015008], including the hierarchical equations of motion (HEOM) [Tan89101, YAN2004216, Jin08234703, PhysRevLett.109.266403], have been developed to simulate non-Markovian open quantum dynamics. However, their applicability remains significantly constrained by the exponential wall inherent to the quantum many-body problem [kohn1999nobel], as their computational costs typically scale exponentially with both system size and environmental memory complexity. To address this challenge, tensor network states (TNSs) [A.Klumper_1993, flannigan2022many, jaschke2018one, link2024open, werner2016positive, Sta183721, Shi18174102, Ke22194102, schwarz2018nonequilibrium, PhysRevLett.123.240602, 10.1063/5.0140002] have been utilized to compress the representation of quantum correlations. However, while TNSs successfully reduce the required number of dynamical variables, the resulting approaches still encounter difficulties when area laws are violated [eisert2010colloquium]. Particularly, TNSs often require large bond dimensions when representing intermediate states during temporal evolution [bonnes2014dynamical, weimer2021simulation] and strong system-environment correlation [Shi18174102].
Neural quantum states (NQSs) [doi:10.1126/science.aag2302], which leverage the power of artificial neural networks, have emerged as a promising method for tackling the exponential wall problem in many-body OQSs [Hermann2023, sharir2020deep, shang2023solving, PhysRevLett.120.240503, donatella2023dynamics, ye2025simulating]. NQSs have demonstrated more expressive power than TNSs [Che18085104, sharir2022neural, lange2024architectures], e.g., in determining the ground states [wu2023tensor, deng2017quantum] and time evolutions [doi:10.1126/science.aag2302] of quantum spin systems. While NQSs have been applied to open quantum systems, simulations have so far been restricted to Markovian environments [PhysRevLett.122.250501, PhysRevLett.122.250502, PhysRevLett.122.250503, mellak2023quantum, PhysRevB.99.214306, mellak2024deep]. Consequently, it is believed that addressing non-Markovian memory effects poses the “ultimate toughness test” for simulating open quantum dynamics [Schuld19]. Moreover, the sign problem inherent to fermionic environments [PhysRevLett.123.050601] introduces an additional challenge [RevModPhys.91.045002].
A recently developed theory [li2024toward] based on the concept of dissipatons (Brownian quasiparticles with complex energies) [Yan14054105] provides an exact description of non-Markovian open quantum dynamics via the dissipaton-embedded quantum master equation (DQME). In this framework, environmental memory is fully encoded in the dissipatons’ statistical behavior. Formulated with second quantization, the DQME treats system and environmental degrees of freedom on an equal footing [li2024toward], making it amenable to a NQS representation of quantum correlations in many-body OQSs.
This Letter aims to establish a framework to integrate the NQS approach with the DQME, and to demonstrate that the unified NQS-DQME approach has the potential to resolve the “ultimate toughness test” of simulating non-Markovian dynamics in strongly correlated OQSs.
Figure 1: Schematic of the fermionic DQME theory, mapping the original OQS (left) to a dissipaton-embedded system (right). Red and blue bars represent the $N_{{}_{\rm S}}$ system fermion energy levels and $N_{{}_{\rm E}}$ memory-carrying dissipaton levels, respectively. Broadening of blue bars indicates each dissipaton’s decay rate (inverse lifetime).
Fermionic DQME method. The total Hamiltonian of an OQS comprises three components: $H_{{}_{\rm T}}=H_{{}_{\rm S}}+H_{{}_{\rm E}}+H_{{}_{\rm SE}}$, where $H_{{}_{\rm S}}$, $H_{{}_{\rm E}}$, and $H_{{}_{\rm SE}}$ represent the system, environment, and system-environment coupling Hamiltonians, respectively. For a noninteracting fermionic environment linearly coupled to the system, the environmental influence is fully characterized by the hybridization time-correlation function. This function is decomposed into a sum of exponentials, $C^{\sigma}(t)=\sum_{j}\eta_{j}^{\sigma}e^{-\gamma_{j}^{\sigma}t}$, which defines the dissipatons [hu2010communication, Yan14054105, li2024toward]. Here, $j$ labels the dissipaton levels, and $\sigma$ denotes the dissipaton charge ($\sigma=+$ for hole-type; $\sigma=$- for electron-type). The imaginary and real parts of $\gamma_{j}^{\sigma}$ correspond to the dissipaton energy and inverse lifetime, respectively, while $\eta^{\sigma}_{j}$ quantifies the system-dissipaton coupling strength. Crucially, the long-time behavior of $C^{\sigma}(t)$, and hence the non-Markovian effects, is dominated by the long-lived dissipatons that are strongly coupled to the system.
Figure 1 illustrates the fermionic DQME theory, where the original OQS is mapped to a dissipaton-embedded system described in terms of the reduced density tensor (RDT), $\bm{\rho}=\rho(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$. Here, vectors $\vec{n}$ and $\vec{n}^{\prime}$ describe system fermion configurations, while $\vec{m}^{-}$ and $\vec{m}^{+}$ represent electron-type and hole-type dissipaton configurations, respectively. The RDT elements with ($\vec{m}^{-},\vec{m}^{+})=\vec{0}$ yield the system’s reduced density operator (RDO). As shown in Fig. 1, each level can accommodate at most one fermion or dissipaton in accordance with the Pauli exclusion principle. In practice, a maximal dissipaton occupation number $M_{\rm max}$ is introduce to facilitate the numerical solution of the DQME [li2024toward].
The fermionic DQME is given by (with $e=\hbar=1$) [li2024toward]:
$\dot{\bm{\rho}}=\mathcal{L}\bm{\rho}=$ $-i[H_{{}_{\rm S}},\bm{\rho}]-\sum_{j}\left(\gamma_{j}^{-}\hat{N}_{j}\bm{\rho}+\gamma_{j}^{+}\bm{\rho}\hat{N}_{j}\right)$
$-i\sum_{j}\Big[\big(\hat{c}_{\nu}^{\dagger}\hat{b}_{j}\bm{\rho}-\hat{b}_{j}\bm{\rho}\hat{c}_{\nu}^{\dagger}\big)+\big(\hat{c}_{\nu}\bm{\rho}\hat{b}^{\dagger}_{j}-\bm{\rho}\hat{b}_{j}^{\dagger}\hat{c}_{\nu}\big)\Big]$
$-i\sum_{j}\left[-\eta_{j}^{-}\,\hat{c}_{\nu}\hat{b}_{j}^{\dagger}\,\bm{\rho}-(\eta_{j}^{+})^{\ast}\,\hat{b}_{j}^{\dagger}\bm{\rho}\hat{c}_{\nu}\right]$
$-i\sum_{j}\left[\eta_{j}^{+}\,\hat{c}_{\nu}^{\dagger}\bm{\rho}\,\hat{b}_{j}+(\eta_{j}^{-})^{\ast}\bm{\rho}\,\hat{b}_{j}\hat{c}_{\nu}^{\dagger}\right].$ (1)
Here, $\hat{c}_{\nu}^{\dagger}$ ($\hat{c}_{\nu}$) creates (annihilates) a fermion at the $\nu$-th system fermion level, $\hat{b}_{j}^{\dagger}$ ($\hat{b}_{j}$) creates (annihilates) a dissipaton at $j$-th level, and $\hat{N}_{j}=\hat{b}_{j}^{\dagger}\hat{b}_{j}$ gives the dissipaton occupation. The RDO of the system, $\rho_{0}$, is obtained by projecting $\bm{\rho}$ onto the dissipaton vacuum.
Fundamentally, dissipatons share similarities with previously proposed pseudomodes [PhysRevA.55.2290, Tam18030402, Lam193721, cirio2023pseudofermion, Lin251289, PhysRevResearch.2.043058], and the DQME is formally equivalent to the exact HEOM theory [Jin08234703, PhysRevLett.109.266403] for reduced system dynamics.
Figure 2: Structure of the neural network representing $\rho_{\rm pre}(\vec{n},\vec{n}^{\prime};\vec{m})$. Boxes group nodes of the same type. The visible layer comprises $2N_{{}_{\rm S}}+N_{{}_{\rm E}}$ nodes, representing elements of $\vec{n}$, $\vec{n}^{\prime}$, and $\vec{m}$, respectively. Connecting lines between boxes indicate that every node in one box is linked to all nodes in the other box through weighted connections.
NQS representation of RDT. We now establish a NQS framework to characterize quantum correlations in OQSs. This framework begins by constructing a preliminary RDT as
$\rho_{\rm pre}(\vec{n},\vec{n}^{\prime};\vec{m})=\sum_{\{\vec{a}\}}\psi(\vec{n};\vec{m};\vec{a})\,\varphi^{\ast}(\vec{n}^{\prime};\vec{m};\vec{a}),$ (2)
where $\vec{m}\equiv(\vec{m}^{-},\vec{m}^{+})$. The wavefunctions $\psi$ and $\varphi$ are expressed using restricted Boltzmann machines (RBMs) [PhysRevB.99.214306, PhysRevLett.122.250501, PhysRevLett.122.250502, PhysRevLett.122.250503, doi:10.1126/science.aag2302], with $\vec{a}$ denoting $N_{\rm a}$ auxiliary nodes of binary values. This construction faithfully captures the mapping illustrated in Fig. 1: the dissipaton states $\vec{m}$ encode the non-Markovian memory of the environment, while $\vec{a}$ accounts for the Markovian background responsible for the finite lifetimes of the dissipatons. Furthermore, this architecture inherently helps preserve the hermiticity and positive definiteness of the resulting RDO.
Specifically,
$\psi(\vec{n};\vec{m};\vec{a})=\sum_{\{\vec{h}\}}{\rm exp}[-E_{\vec{h}}(\vec{n};\vec{m};\vec{a})],$ (3)
where $\vec{h}$ represents $N_{\rm h}$ hidden nodes, and
$E_{\vec{h}}(\vec{n};\vec{m};\vec{a})=\bm{a}^{T}\bm{b}+\bm{n}^{T}\bm{c}+\bm{m}^{T}\bm{d}+\bm{h}^{T}\bm{g}+\bm{n}^{T}\bm{D}\bm{m}$
$\qquad\quad+\bm{m}^{T}\bm{K}\bm{a}+\bm{n}^{T}\bm{X}\bm{h}+\bm{n}^{T}\bm{X}^{\prime}\bm{a}+\bm{m}^{T}\bm{Y}\bm{h}.$ (4)
Here, $(\bm{b},\bm{c},\bm{d},\bm{g})$ denote complex-valued bias vectors, while $(\bm{D},\bm{K},\bm{X},\bm{X}^{\prime},\bm{Y})$ represent complex-valued weight matrices of the RBM. An analogous RBM structure describes $\varphi(\vec{n}^{\prime};\vec{m};\vec{a})$.
Figure 2 depicts the neural network architecture for $\rho_{\rm pre}(\vec{n},\vec{n}^{\prime};\vec{m})$. The complete analytic expression is provided in the Supplemental Material (SM) [SM]. The dissipaton-related nodes in the visible layer explicitly encode the environmental memory content, in distinct contrast to conventional RBMs used to represent OQSs coupled to memoryless environments [PhysRevLett.122.250501, PhysRevLett.122.250502, PhysRevLett.122.250503].
The dimensionality of RDT can be reduced by leveraging its inherent symmetry and sparsity [10.1063/1.4914514]. First, we enforce the symmetry constraint by constructing
$\rho_{\rm sym}(\vec{s})=\rho_{\rm pre}(\vec{s})+(-1)^{\lfloor M^{-}/2\rfloor+\lfloor M^{+}/2\rfloor}\rho_{\rm pre}^{\ast}(\vec{s}^{\,T}),$ (5)
where $\lfloor\cdot\rfloor$ denotes the floor function, $M^{\sigma}=\sum_{j}m_{j}^{\sigma}$, and $m_{j}^{\sigma}$ is the $j$-th element of $\vec{m}^{\sigma}$. Here, $\vec{s}\equiv(\vec{n},\vec{n}^{\prime};\vec{m}^{-},\vec{m}^{+})$ represents a visible state, while $\vec{s}^{\,T}\equiv(\vec{n}^{\prime},\vec{n};\vec{m}^{+},\vec{m}^{-})$ is its block-swapped counterpart. Next, we introduce a filter function $f_{\rm{spa}}(\vec{s})$ that identifies, a priori, the zero elements of $\bm{\rho}$ based on the sparsity pattern dictated by the forms of $H_{{}_{\rm S}}$ and $H_{{}_{\rm SE}}$ [Hou15104112]. In practice, $f_{\rm{spa}}(\vec{s})$ normally reduces the $\vec{s}$-space by over an order of magnitude [SM]. The resulting symmetrized and filtered RDT is given by $\rho(\vec{s})=f_{\rm spa}(\vec{s})\rho_{\rm sym}(\vec{s})$.
Within the RBM representation, the RDT is expressed parametrically as $\bm{\rho}_{\bm{\alpha}}$, where $\bm{\alpha}$ denotes the complete set of RBM parameters. The total number of parameters is approximately $N_{\rm para}\approx 2N_{{}_{\rm S}}(N_{\rm h}+N_{\rm a})+N_{{}_{\rm E}}(4N_{\rm h}+N_{\rm a})$. Provided that $N_{\rm h}$ and $N_{\rm a}$ scale linearly with $2N_{{}_{\rm S}}+N_{{}_{\rm E}}$, $N_{\rm para}$ scales approximately quadratically with the dimension of the dissipaton-embedded system [SM]. This scaling is vastly more favorable than that of the original DQME, where the RDT dimension grows as $N_{\rm RDT}\approx 4^{N_{{}_{\rm S}}}N_{{}_{\rm E}}^{M_{\rm max}}$.
Time evolution of NQS-DQME. Within the NQS framework, the exact DQME of Eq. (1) is projected onto the parameter space. Through the time-dependent variational principle [PhysRevLett.122.250502, reh2021time], the problem is transformed into determining the time evolution of the parameters $\bm{\alpha}(t)$. Numerically, this is achieved by minimizing the 2-norm $\Delta{s}^{2}=\|\dot{\bm{\rho}}_{\bm{\alpha}}-\mathcal{L}{\bm{\rho}}_{\bm{\alpha}}\|^{2}_{2}$, which yields a linear equation for $\dot{\bm{\alpha}}$ at each time step:
$\mathcal{S}\dot{\bm{\alpha}}=\mathcal{F}.$ (6)
Here, $\mathcal{S}$ and $\mathcal{F}$ are a matrix and a vector defined in the parameter space, respectively. Evaluating their elements requires summming over all visible states $\vec{s}$, which we perform using the Markov chain Monte Carlo algorithm [geyer1992practical]. The sampling process employs the Metropolis-Hastings criterion [metropolis1953equation] with a customized probability distribution $p(\vec{s})\propto e^{-3(M^{+}+M^{-})}$. The derivatives $\dot{\bm{\alpha}}$ are then obtained by solving Eq. (6) [minres], and $\bm{\alpha}(t)$ are updated using the fourth-order Runge-Kutta algorithm [press2007numerical, caoge2024code]. Further details are provided in the SM [SM].
To demonstrate the practicality and accuracy of the NQS-DQME approach, we investigate non-Markovian dynamics in two strongly correlated OQSs. We employ the numerically exact fermionic HEOM method implemented in the HEOM-QUICK2 program [zhang2024heom] as the benchmark to provide reference values.
Figure 3: (a) Schematic of open quantum dynamics for an impurity coupled to left ($L$) and right ($R$) reservoirs with chemical potentials $\mu_{L}$ and $\mu_{R}$. Gray shaded regions in the reservoirs represent Kondo clouds screening the impurity’s localized spin. (b) $I_{R}(t)$ at different temperatures. The NQS-DQME results (solid lines, vertically offset for clarity) are benchmarked against HEOM reference values. Inset: steady-state current versus inverse temperature. (c) Time evolution of $\Delta s^{2}$ for $k_{\rm B}T=0.3\,\Gamma$, computed with different $N_{\rm{h}}=N_{\rm{a}}$. Inset: $\mathcal{E}_{I_{R}}$ versus $N_{\rm{h}}$ with a power function fit. (d) $I_{R}(t)$ at $k_{\rm B}T=0.3\,\Gamma$ computed with different $N_{\rm MC}$. Inset: $\mathcal{E}_{I_{R}}$ versus $N_{\rm MC}$. (e) Number of dynamical variables explicitly accessed in the NQS-DQME and HEOM methods versus $N_{{}_{\rm E}}$, where $N_{{}_{\rm E}}$ increases monotonically as $T$ decreases. Dashed lines are power function fits to the scattered data with $M_{\rm max}=3$. System parameters (in units of $\Gamma$): $\epsilon_{0}=U_{0}/2=2$, $\Delta\epsilon=-7$, and $\Delta U=6$; see the SM for further details [SM].
Case 1: non-Markovian charge relaxation with emergent Kondo correlations. We consider an OQS comprising a localized impurity symmetrically coupled to two noninteracting electron reservoirs. This setup is important for understanding electron transport through molecular junctions. The OQS is described by the single-impurity Anderson model [And6141]. The impurity Hamiltonian is $H_{{}_{\rm S}}(t)=\epsilon_{0}(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+U_{0}\hat{n}_{\uparrow}\hat{n}_{\downarrow}+{\Theta}(t-t_{0})[\Delta\epsilon(\hat{n}_{\uparrow}+\hat{n}_{\downarrow})+\Delta U\hat{n}_{\uparrow}\hat{n}_{\downarrow}$], where $\hat{n}_{s}$ is the occupation operator for spin-$s$ electrons, $\epsilon_{0}$ is the impurity energy, and $U_{0}$ is the Coulomb interaction energy. At time $t_{0}=0$, a sudden quench shifts the impurity’s parameters by $\Delta\epsilon$ and $\Delta U$, and a bias voltage is applied, establishing a chemical potential difference between the reservoirs. Figure 3(a) depicts the ensuing open quantum dynamics. The impurity level shift triggers electron transfer from the reservoirs to the impurity on a relatively short timescale. Subsequently, reservoir electrons redistribute their spins to screen the impurity’s localized spin, leading to the formation of Kondo states at the impurity-reservoir interfaces over a longer timescale [anders2005real, PhysRevLett.119.156601, Din24174120]. We denote $t_{\rm sh}$ and $t_{\rm lo}$ as representative times for the short-time and long-time regions, respectively.
Figure 3(b) shows the time-dependent current $I_{R}(t)$ at various temperatures. At low temperatures, $I_{R}(t)$ exhibits pronounced oscillations near $t=1/\Gamma$, where $\Gamma$ is the impurity-reservoir hybridization strength. These transient oscillations are a characteristic signature of non-Markovian memory, arising from coherent impurity-reservoir coupling under suppressed thermal fluctuations. The same coupling enhances Kondo correlations, leading to the increase in the steady-state current (evaluated at $t_{\rm lo}$) as temperature decreases, as seen in the inset.
Within the NQS-DQME framework, the expectation value of an observable $\hat{X}$, $X=\langle\hat{X}\rangle$, is evaluated from $\bm{\rho}_{\bm{\alpha}}$ [SM]. The accuracy of the results is quantified by the relative error:
$\mathcal{E}_{X}\equiv\frac{\int_{t_{0}}^{t_{{\rm lo}}}\lvert X(t)-X^{{{\rm ref}}}(t)\rvert\,{\rm d}t}{\int_{t_{0}}^{t_{{\rm lo}}}\frac{1}{2}\big(\lvert X(t)\rvert+X^{{{\rm ref}}}(t)\rvert\big){\rm d}t}.$ (7)
As shown in Fig. 3(b), the NQS-DQME results exhibit excellent agreement with the reference data, with $\mathcal{E}_{I_{R}}<1\%$ across the entire temperature range.
Figure 4: Visualization of the RBM parameters for the open quantum dynamics shown in Fig. 3. (a,b) Heatmaps of $\lvert\bm{K}(t_{\rm lo})-\bm{K}(t_{\rm sh})\rvert$ at (a) $k_{\rm B}T=3.0\,\Gamma$ and (b) $k_{\rm B}T=0.3\,\Gamma$, for $t_{{\rm sh}}=1.25/\Gamma$ and $t_{{\rm lo}}=5/\Gamma$. The vertical axis indexes dissipaton levels (with decay rates sorted from slowest to fastest), and the horizontal axis indexes the auxiliary nodes. (c,d) Time evolution of the auxiliary-layer-averaged weight difference for the $j$th dissipaton level, $\Delta K_{j}=\lvert K_{j}(t)-K_{j}(t_{\rm sh})\rvert$, at (c) $k_{\rm B}T=3.0\,\Gamma$ and (d) $k_{\rm B}T=0.3\,\Gamma$. Here, $\gamma_{j}\equiv{\rm Re}(\gamma^{\sigma}_{j})$ denotes the decay rate in unit of $1/\Gamma$. The red line tracks the slowest-decaying dissipaton level.
To explore the tradeoff between numerical cost and accuracy, we examine the low temperature regime ($k_{\rm B}T=0.3\,\Gamma$), where non-Markovian effects are prominent. The dissipaton-embedded system comprises $2$ system levels coupled inhomogeneously to $64$ dissipaton levels. Figure 3(c) shows the evolution of the precision metric $\Delta s^{2}$ for different numbers of hidden nodes $N_{\rm h}$. The magnitude of $\Delta s^{2}$ decreases by orders of magnitude with a linear increase in $N_{\rm h}$, demonstrating rapid convergence. Similarly, Fig. 3(d) shows that that accurate results are achieved with relatively few Monte Carlo samples ($N_{\rm MC}$). Furthermore, Fig. 3(e) demonstrates that the NQS-DQME approach achieves high accuracy ($\Delta s^{2}<10^{-8}$) while utilizing drastically fewer dynamical variables than the conventional HEOM method. This comparison is conducted under equivalent conditions, as the same symmetry and sparsity relations are incorporated into both methods. These results collectively highlight the outstanding expressive power of the RBM for characterizing non-Markovian open quantum dynamics.
An important hypothesis in open quantum dynamics is that non-Markovian memory is crucial for long-time phenomena like the formation of Kondo-correlated states. However, directly verifying this by tracking dynamical variables in conventional methods such as the HEOM or the original DQME is prohibitively difficult due to their high dimensionality. In contrast, the NQS-DQME framework makes such tracking feasible, since the RDT is represented by weight matrices that form a much lower-dimensional parameter space.
This capability is demonstrated in Fig. 4, which visualizes the evolution of the RBM weight matrix $\bm{K}$ connecting dissipaton-level nodes to auxiliary nodes for the open quantum dynamics shown in Fig. 3. A comparison of the heatmaps in Fig. 4(a) and (b) shows that the weight changes, $\lvert\bm{K}(t_{\rm lo})-\bm{K}(t_{\rm sh})\rvert$, are significantly more pronounced for the slowest-decaying dissipaton level at the lower temperature. This is indicated by the darker top row in Fig. 4(b). This observation is further confirmed by the time evolution of the auxiliary-layer-averaged weight difference $\Delta K_{j}(t)$ in Fig. 4(c) and (d), where the slowest-decaying dissipaton level (red line) becomes increasingly prominent as temperature decreases, directly linking the non-Markovian memory to the emerging Kondo effect.
Figure 5: (a) Schematic of quantum dissipative dynamics for two impurities coupled to a reservoir. Spin-exchange interaction (wavy line) and Kondo clouds (gray regions) are shown, where cloud extents indicate Kondo correlation strengths. Time evolution of (b) $\dot{S}_{12}$ and (c) $E_{\rm hyb}$ at various temperatures. NQS-DQME results (solid lines, vertically offset for clarity) are benchmarked against HEOM reference values. Inset in (c): $\Delta E_{\rm hyb}=E_{\rm hyb}(t_{\rm lo})-E_{\rm hyb}(t_{0})$ versus inverse temperature. The increase in $\Delta E_{\rm hyb}$ with decreasing temperature indicates weakened Kondo correlations in the long-time limit. System parameters (in units of $\Gamma$): $\epsilon_{0}=-U_{0}/2=-6$, and $J=8$; see the SM for details [SM].
Case 2: non-Markovian spin relaxation with intertwined Kondo correlations. We now examine a more complex OQS comprising two impurities coupled to a reservoir. This setup directly models recent experimental studies of coherent spin dynamics in surface molecules [doi:10.1126/science.aay6779, doi:10.1126/science.abg8223]. The system Hamiltonian is $H_{{}_{\rm S}}(t)=\sum_{\nu=1,2}[\epsilon_{0}(\hat{n}_{\nu\uparrow}+\hat{n}_{\nu\downarrow})+U_{0}\hat{n}_{\nu\uparrow}\hat{n}_{\nu\downarrow}]-\Theta(t-t_{0})J\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2},$ where the final term activates a ferromagnetic exchange interaction ($J>0$) between the impurities at time $t_{0}$.
Figure 5(a) shows the relaxation dynamics following the sudden activation of the exchange interaction. The initially independent localized spins on the two impurities gradually evolve towards parallel alignment. This results in a localized high-spin state that cannot be fully screened by the reservoir electrons. This underscreening progressively suppresses Kondo correlations during the temporal evolution [PhysRevLett.103.197202, parks2010mechanical, li2020molecular].
Figure 5(b) and (c) depict the time derivative of the inter-impurity spin correlation, $\dot{S}_{12}=\frac{\partial}{\partial t}\langle\hat{\bm{S}}_{1}\cdot\hat{\bm{S}}_{2}\rangle$, and the impurity-reservoir hybridization energy, $E_{\rm hyb}=\langle H_{{}_{\rm SE}}\rangle$, respectively, at various temperatures. We plot $\dot{S}_{12}$ rather than ${S}_{12}$ to highlight its more pronounced oscillatory behavior. At low temperatures, both quantities exhibit strong oscillations, a hallmark of non-Markovian memory effects.
The NQS-DQME results are in excellent agreement with the reference values, with maximum errors $\mathcal{E}_{{S}_{12}}$ and $\mathcal{E}_{{E}_{\rm hyb}}$ both below $1\%$ for $t_{\rm lo}>3/\Gamma$. This quantitative agreement is remarkable, as the spin relaxation dynamics are governed by the complex interplay of direct inter-impurity correlations and impurity-reservoir Kondo correlations. Furthermore, similar to Case 1, the NQS representation achieves this accuracy with a significantly reduced number of dynamical variables compared to conventional approaches [SM].
Summary. We have developed a NQS framework for the efficient simulation of non-Markovian open quantum dynamics by integrating the dissipaton concept into the ansatz. This NQS-DQME approach provides a compact and physically interpretable epresentation of the RDT, enabling the simultaneous treatment of complex environmental memory and many-body correlations with significantly enhanced computational efficiency.
While the current framework shares a common challenge with many NQS methods, i.e., the computationally demanding minimization of the loss function $\Delta s^{2}$ [PhysRevLett.122.250503]. This bottleneck can be addressed using advanced optimizers and more sophisticated network architectures. Given its proven scalability and interpretability, the NQS-DQME approach presents a promising and versatile platform for exploring rich quantum phenomena in correlated, non-Markovian settings.
Acknowledgements.
Support from the National Natural Science Foundation of China (Grants No. 22393912, No. 22321003, No. 22103073, No. 22173088, and No. 22373091), the Innovation Program for Quantum Science and Technology (Grants No. 2021ZD0303301 and No. 2021ZD0303306), the AI for Science Foundation of Fudan University (Grant No. FudanX24AI023), and the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB0450101) is gratefully acknowledged.